OpenStudy (anonymous):

Using the given zero, find all other zeros of f(x). -2i is a zero of f(x) = x^4 - 21x^2 - 100

OpenStudy (anonymous):

Here are the choices (: 2i, 10, -10 2i, 5, -5 2i, 10i, -10i 2i, 5i, -5i

OpenStudy (amoodarya):

if all coefficent are real numbers then if a+ib is root a-ib is also root of equation

OpenStudy (amoodarya):

so another root is +2i divide f(x) to (x+2i)(x-2i)

OpenStudy (amoodarya):

another solution is to use "a=x^2" so f=a^2-21a-100

OpenStudy (anonymous):

So, would the answer be 2i, 5, -5???

OpenStudy (amoodarya):

yes

OpenStudy (anonymous):

Thank you so much!! :D